OpenStudy (anonymous):

Charlotte has two job offers as a car sales person. The first job will pay her $600 a month plus 2% on the price of her car sales. The second job will pay $1000 a month plus 1% on the price of her sales. How much must she sell each month to make the same income from the two jobs? A. $20,000 B. $30,000 C. $40,000 D. $50,000

hero (hero):

Proper Solution: If we Let j1 = 1st job pay j2 = 2nd job pay Then: j1 = 600 + .02x j2 = 1000 + .01x The question asks, at what income will the 1st job pay equal the 2nd job pay. In other words, when will j1 = j2 ? To find this set the expressions for j1 and j2 equal to each other, then solve for x as follows: j1 = j2 600 + .02x = 1000 + .01x .02x - .01x = 1000 - 600 .01x = 400 x = 400/.01 x = 40,000

hero (hero):

Actually, the way that they ask the question is a bit misleading and it shouldn't have been asked that way. I can see how you were confused about it.

OpenStudy (anonymous):

The key is equate the two jobs: $600 + 2% S = $ 1,000 + 1% S